# Adventures in Algebraic Geometry

### v1

@@
For a nonzero $f(z)$, its zeros and poles come in pairs $(a,-a)$ modulo $L$. Moreover, the multiplicity of zero/pole at $a$ with $2a\in L$ is even, see Exercise 2. Then we consider
@@
Exercise 1. Prove that for every lattice $L$ there exists a constant $c$ such that the number of $l\in L$ with $k\leq|l|<k+1$ is at most $c\,k$ for all $k\geq 1$. Hint: Argue that the union of fundamental parallelograms of $L$ centered at such $l$ lies in a certain annulus and thus has smaller area.
Exercise 2. Prove that for any even elliptic function $f$ and any $a\in\frac{1}{2}L$ the order of pole or zero of $f$ at $a$ is even. Hint: Consider the Laurent power series of $f$ at $a$ and use $f(2a-z)=f(-z)=f(z)$.
@@
and the elliptic functions $\frac{\mathcal{P}(z)}{\mathcal{P}^{\prime}(z)}$ and $\frac{1}{\mathcal{P}^{\prime}(z)}$ are holomorphic near $z=0$ and have limit 00 as $z\to 0$. ∎
@@
Since the three poles of this function are all at $0\,{\rm mod}\,L$, the sum of these zero is $0\,{\rm mod}\,L$ by Proposition 11.4. Similarly, for $\alpha_{1}=0$, but $a_{0}\neq 0$, the line intersect $\mu({\mathbb{C}}/L)$ at $\mu(0)$ and two other points that add up to zero. If we start with $A=\mu(z_{1})$, $B=\mu(z_{2})$, then the third intersection point of $AB$ with $\mu({\mathbb{C}}/L)$ is $\mu(-z_{1}-z_{2})$, and then the point $A+B$ corresponds to $\mu(z_{1}+z_{2})$, so the two group laws are matched under $\mu$12.212.212.2 We are ignoring various issues of multiple roots, or one of $A$ or $B$ or $A+B$ being equal to $(0:1:0)$. These can be handled by continuity, or directly.
@@
Uniqueness is clear because the ratio of two functions with the same multiplicities of zeros and poles is holomorphic on ${\mathbb{C}}/L$ and thus constant. Existence is harder to prove – we will do it with the help of the Jacobi theta function which is a holomorphic function on ${\mathbb{C}}\times\mathcal{H}$, where =$\mathcal{H}=\{\tau,{\rm~{}Im}\tau>0\}$ is the upper half plane.
@@
We lift $z_{k}$ to ${\mathbb{C}}$. We would like to be able to do it so that $\sum_{k}m_{k}z_{k}=0$, but a priori we can only get $\sum_{k}m_{k}z_{k}\in L$, if none of the $m_{k}$ is equal to ±$\pm 1$. However, we can also split some $m_{k}$ into $m_{k}-1$ and $1$ and do a separate lift for $1$ to get the sum to be exactly zero. Now the elliptic function $f$ from Proposition 13.3 defines a meromorphic function on ${\mathbb{C}}/L$ with prescribed zeros and poles. (Note that we did not assume in Proposition 13.3 that $z_{k}$ are pairwise distinct modulo $L$.) ∎
@@
The torsion is well-understood, but the rank is harder. For example, it is unknown whether the rank $n$ is bounded from above. The current world record of $n=28$ was achieved by Elkies in 2006, see [13].
@@
$-{\mathrm{i}}\sum_{n\in{\mathbb{Z}}}(-1)^{n}{\mathrm{e}}^{\pi{\mathrm{i}}((n+\frac{1}{2})^{2}\tau+(2n+1)z)}$
@@
First, we note that in the case of $X={\mathbb{C}}{\mathbb{P}}^{1}$ the degree of $K_{X_{{\mathbb{C}}{\mathbb{P}}^{1}}}$ is $(-2)$. Indeed, if we look at the differential form $dz$ which has no zeros/poles in ${\mathbb{C}}$, it has a second order pole at infinity, because $dz=-\left(\frac{1}{z}\right)^{-2}\,d\left(\frac{1}{z}\right)$ near $\infty$.
@@
For $g(X)=1$, the automorphism group of $X={\mathbb{C}}/L$ is a semidirect product of ${\mathbb{C}}/L$ with a finite group, which could be ${\mathbb{Z}}/2{\mathbb{Z}}$, ${\mathbb{Z}}/4{\mathbb{Z}}$ or ${\mathbb{Z}}/6{\mathbb{Z}}$. The normal subgroup ${\mathbb{C}}/L$ is obtained by parallel transports $\varphi_{w}:z\mapsto z+w$, and the finite group comes from scalar multiplications that preserve the lattice $L$. It is usually just {±$\{\pm 1\}$, but it is $\{1,{\mathrm{i}},-1,{\mathrm{i}}\}$ if $L={\mathbb{Z}}+{\mathbb{Z}}{\mathrm{i}}$ and it is the group of sixth roots of $1$ for the hexagonal lattice $L={\mathbb{Z}}+{\mathbb{Z}}(\frac{1}{2}+\frac{\sqrt{3}}{2}{\mathrm{i}})$. We do not prove this, but it is not too hard to derive from what we know about elliptic curves.
@@
$-2+\frac{1}{2}+\frac{2}{3}+\frac{6}{7},$
@@
[…]
@@
Exercise 3. Prove the last claim of the section. Hint: First argue that for $k\geq 5$ the right hand side is at least $\frac{1}{2}$. Then for $k=4$ the right hand is positive, so at least one of $n_{i}$ is $3$ or more, which gives at least $-2+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{2}{3}=\frac{1}{6}$. If $k\leq 2$, the right hand side is negative, so we must have $k=3$, etc.
@@
$S=\{(x_{1},\ldots,x_{n})\times(y_{1}:\ldots:y_{n}),{\rm~{}such~{}that~{}rk}\hskip-2.0pt\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{n}\\ y_{1}&\cdots&y_{n}\end{array}\right)=1\}.$
@@
    $X=X_{0}\dashleftarrow\hskip-5.0pt\dashrightarrow X_{1}\dashleftarrow\hskip-5.0pt\dashrightarrow\cdots\dashleftarrow\hskip-5.0pt\dashrightarrow X_{n-1}\dashleftarrow\hskip-5.0pt\dashrightarrow X_{n}=Y$
@@
Now that we have established that $({\mathbb{C}}[x_{1},\ldots,x_{n}]/I)^{G}$ is finitely generated, let us prove the bijection between its maximal ideals and the set of orbits of $G$-action on the set of maximal ideals of ${\mathbb{C}}[x_{1},\ldots,x_{n}]/I$16.116.116.1 Recall that points in $X$ correspond to maximal ideals in $R={\mathbb{C}}[x_{1},\ldots,x_{n}]/I$ by Hilbert’s Nullestellensatz. In one direction, to any maximal ideal $m\subset R$ we associate the ideal $m^{G}=m\cap R^{G}\subset R^{G}$. The ideal $m^{G}$ is maximal because for any $x\in R^{G}\setminus m^{G}$ there exists $y\in R$ such that $xy=1\,{\rm mod}\,m$. Then the element
@@
on it. This gives a group homomorphism ${\rm SU(2)}\to{\rm SO}(3,{\mathbb{R}})$, whose kernel is {±$\{\pm 1\}$. Finite subgroups of ${\rm SO}(3,{\mathbb{R}})$ are classically known. They include two infinite series (rotation symmetries of a regular pyramid and a regular bipyramid) as well as rotational symmetries of the regular tetrahedron, cube/octahedron, and icosahedron/dodecahedron. This allows one to classify all finite subgroups $G\subset{\rm SL}(2,{\mathbb{C}})$.
@@
[…]
Let us see this with all the gory details, in the coordinate charts. Recall that the blowup ~$\tilde{\mathbb{C}}^{3}$ is a submanifold of ${\mathbb{C}}^{3}\times{\mathbb{C}}{\mathbb{P}}^{2}$, given by $\{(u,v,w),(u_{1}:v_{1}:w_{1})\}$ such that there exists $\lambda$ with $(u,v,w)=\lambda(u_{1},v_{1},w_{1})$. As a result, ~$\tilde{\mathbb{C}}^{3}$ is covered by three open charts, each one isomorphic to ${\mathbb{C}}^{3}$.
@@
We thus see that the proper preimage $\tilde{S}$ of $S$ is smooth. It is also instructive to look at the fibers of $\tilde{S}\to S$. Outside of $(0,0,0)$, the fibers are of course points. The preimage of $(0,0,0)$ is a subvariety of the exceptional divisor, given by $(u_{1}:v_{1}:w_{1})\in{\mathbb{C}}{\mathbb{P}}^{2}$ with $u_{1}v_{1}-w_{1}^{2}=0$, so it is a ${\mathbb{C}}{\mathbb{P}}^{1}$. For example, in the coordinate chart $w_{1}\neq 0$, we have $\frac{u}{w}=\frac{u_{1}}{w_{1}},~{}\frac{v}{w}=\frac{v_{1}}{w_{1}}$ so $10=\left(\frac{u}{w}\right)\left(\frac{v}{w}\right)-$ 1 matches $u_{1}v_{1}-w_{1}^{2}=0$ in the appropriate chart on ${\mathbb{C}}{\mathbb{P}}^{2}$.
@@
We repeat the process. Each time, the power $n$ decreases by $2$ and two new ${\mathbb{C}}{\mathbb{P}}^{1}$-s spring out. A paranoid reader can verify that under the blowup the previous ${\mathbb{C}}{\mathbb{P}}^{1}$-s separate and two new lines intersect them transversely. The last step is either unnecessary (if $n$ is odd, and we reduce to $uv=w$) or we only have one ${\mathbb{C}}{\mathbb{P}}^{1}$ (if $n$ is even and we reduce to $A_{1}$-singularity). In both cases we end up with a resolution of singularities $\tilde{S}\to S$ such that the preimage of the singular point is a chain of $(n-1)$ transversal ${\mathbb{C}}{\mathbb{P}}^{1}$-s.
@@
with $n-1$ vertices, which is the $A_{n-1}$ Dynkin diagram.17.317.317.3 There is, of course, more to this notation. Appropriately understood intersection pairing matrix of the components is up to the sign the Cartan matrix of $A_{n-1}$. We will not try to define it now, but will comment on it in Section 19.
@@
The symmetry $(x,y)\mapsto(-y,x)$ now becomes the involution $(u,v,w)\mapsto(v,u,-w)$. We can take the invariants first and then the quotient. The action of this involution on ${\mathbb{C}}^{3}$ has positive eigenspace $u+v$ and the $A_{1}$-type action on $u-v$ and $w$, so the ring of invariants is generated by
@@
$02x=0,~{}z+(n-1)z^{n+1}y^{n-2}=0,~{}y+(n+1)z^{n}y^{n-1}=0,~{}x^{2}+yz+z^{n+1}y^{n-1}=$ 0 .
We can multiply the second equation by $(n+1)y$ and the third one by $(n-1)z$, subtract them and get $yz=0$. Since $n>2$, we use the second and third equations to deduce that both $y$ and $z$ are 00. So the only singular point in this chart is $(0,0,0)$, which corresponds to $(0:1:0)$ on the exceptional ${\mathbb{C}}{\mathbb{P}}^{2}$ of ~$\tilde{\mathbb{C}}^{3}\to{\mathbb{C}}^{3}$ (we are in the chart $v_{1}\neq 0$). For $n>2$, we see that the singularity at this point is of type $A_{1}$, because we can ignore higher degree terms. The intersection with the exceptional locus $v=0$ is given by $\frac{u}{v}=0$, i.e. we have the line (0::)(0:*:*)( 0 : ∗ : ∗ ) on the exceptional ${\mathbb{C}}{\mathbb{P}}^{2}$ of ~$\tilde{\mathbb{C}}^{3}\to{\mathbb{C}}^{3}$.
@@
[…]
@@
and the action of $G/G_{1}\cong{\mathbb{Z}}/2{\mathbb{Z}}$ which switches $u$ and $v$. The resolution $S_{1}$ has a chain of $(2n-1)$ Riemann spheres
@@
which intersect transversely, and the involution of $G/G_{1}$ switches the opposite spheres. Because we have an odd number of spheres, this involution will preserve the middle ${\mathbb{C}}{\mathbb{P}}^{1}$. It acts by switching two intersection points with two adjacent ${\mathbb{C}}{\mathbb{P}}^{1}$-s. We can call these points 00 and $\infty$ and observe that every such involution can be written in the form $z\mapsto\frac{1}{z}$, see Exercise 1. Then we have two fixed points $z=\pm 1$. Since these are isolated points, the eigenvalues of the action on the tangent space must be $(-1,-1)$, so the quotient will have an $A_{1}$ singularity there. Therefore, the image of the chain of $(2n-1)$ lines on $S_{2}$ will be $n$ lines, with the first line having two additional $A_{1}$ singularities on it. Blowing up these $A_{1}$ singularities, we get $S_{3}$ with the $(n+2)$ lines in $D_{n+2}$ configuration.
We can similarly get the $E_{6}$ singularity as the quotient of the $D_{4}$ singularity by a ${\mathbb{Z}}/3{\mathbb{Z}}$ action. Then if you resolve the $D_{4}$ singularity to get a surface with a $D_{4}$ configurations of projective lines, the group can be seen rotating the three “outside” lines while fixing the middle one. By Exercise 2, any order three automorphism of ${\mathbb{C}}{\mathbb{P}}^{1}$ can be written in the form $z\mapsto{\mathrm{e}}^{2\pi{\mathrm{i}}/3}z$, so it will have two fixed points 00 and $\infty$. It is then possible to compute that the quotient will be the $A_{2}$ singularity, so we get $S_{1}\to S$ with $E_{6}$ singularity replaced by two lines with two $A_{2}$ singular points on one of them.
@@
Exercise 1. Prove that every involution of ${\mathbb{C}}{\mathbb{P}}^{1}$ that sends 00 to $\infty$ can be written in the form $z\mapsto\frac{c}{z}$ for some nonzero constant $c$. Then argue that by scaling $z$ we can rewrite it as $z\mapsto\frac{1}{z}$.
@@
If $\,Y$ and $Z$ intersect transversely at $k$ points, then $Y\cdot Z=k$ (in particular, $Y\cdot Z=0$ if $Y\cap Z=\emptyset$). More generally, if $Y\cap Z$ is a finite set, then $Y\cdot Z$ is a sum of local contributions, each of which is positive. However, the definition makes sense even if $Y\cap Z$ is positive-dimensional, or even $Y\subseteq Z$! In these cases, we just know that $Y\cdot Z$ is integer but can make no claims about the sign of it. Philosophically, we can deform $Y$ as a real cycle to some $Y^{\prime}$ with the same homology classes $[Y^{\prime}]=[Y]$ so that $Y^{\prime}$ intersects $Z$ transversely, and we add up (±$(\pm 1)$ at each intersection point, depending on the orientations.
@@
The key observation is that $L\sim L^{\prime}+E$, in the sense that $[L]=[L^{\prime}]+[E]$ in $H_{2}(X,{\mathbb{Z}})$. Indeed, more generally there is a map of Weil divisors to $H_{2}(X,{\mathbb{Z}})$ which passes through the class group. Another way of saying the same thing is that a divisor of a rational function is always homologically trivial. And there is a rational function (the ratio of the equations of the equations of the two lines) which has simple zeros at $E$ and $L^{\prime}$ and poles at $L$19.319.319.3 In this particular case can think of the family of smooth preimages of lines in ${\mathbb{C}}{\mathbb{P}}^{2}$ degenerating into a singular curve $E\cup L^{\prime}$. We also have
@@
$[C_{1}^{\prime}]+[E]+[C_{2}^{\prime}]-[D]$
where $D$ is the preimage of the conic at infinity $\{uv-w^{2}=t=0\}$. Therefore, we have $[E]=[D]-[C_{1}^{\prime}]-[C_{2}^{\prime}]$ in $H_{2}(X,{\mathbb{Z}})$ and
@@
Exercise 3. Let $X$ be a blowup of ${\mathbb{C}}{\mathbb{P}}^{2}$ in $5$ points $p_{1},\ldots,p_{5}$ in general position. Prove that the proper preimage of any line through two of these points has self-intersection $(-1)$. Prove that the proper preimage of the unique conic that passes through all five points also has self-intersection $(-1)$.
@@
One can show that every Weil divisor on ${\mathbb{C}}^{2}$ can be given as a divisor of a rational function. Therefore, the class group of $H_{n}$ is generated by $[S_{0}]$ and $[F_{0}]$. Let us find their intersection numbers.
@@
The rational function $\frac{x_{0}}{x_{1}}$ on ${\mathbb{C}}{\mathbb{P}}^{1}$ gives a rational function on $H_{n}$ with divisor $[F_{0}]-[F_{\infty}]$. We thus get
@@
$[S_{0}]-[S_{-n}]+n[F_{0}].$ (20.2)
This would imply that $[S_{0}]=[S_{-n}]-n[F_{0}]$ in ${\rm Cl}(H_{n})$, so
@@
has divisor $[S_{0}]-[S_{-n}]$ in the first chart and $[S_{0}]-[S_{-n}]+n[F_{0}]$ in the second. ∎

### v2

@@
For a nonzero $f(z)$, its zeros and poles come in pairs (a,a)(a,-a) modulo $L$. Moreover, the multiplicity of zero/pole at $a$ with $2a\in L$ is even, see Exercise 2. Then we consider
@@
Exercise 1. Prove that for every lattice $L$ there exists a constant $c$ such that the number of $l\in L$ with $k\leq|l|<k+1$ is at most ckc\,k for all $k\geq 1$. Hint: Argue that the union of fundamental parallelograms of $L$ centered at such $l$ lies in a certain annulus and thus has smaller area.
Exercise 2. Prove that for any even elliptic function $f$ and any $a\in\frac{1}{2}L$ the order of pole or zero of $f$ at $a$ is even. Hint: Consider the Laurent power series of $f$ at $a$ and use f(2az)=f(z)=f(z)f(2a-z)=f(-z)=f(z).
@@
and the elliptic functions $\frac{\mathcal{P}(z)}{\mathcal{P}^{\prime}(z)}$ and $\frac{1}{\mathcal{P}^{\prime}(z)}$ are holomorphic near $z=0$ and have limit 0 as $z\to 0$. ∎
@@
Since the three poles of this function are all at $0\,{\rm mod}\,L$, the sum of these zero is $0\,{\rm mod}\,L$ by Proposition 11.4. Similarly, for $\alpha_{1}=0$, but $a_{0}\neq 0$, the line intersect $\mu({\mathbb{C}}/L)$ at $\mu(0)$ and two other points that add up to zero. If we start with $A=\mu(z_{1})$, $B=\mu(z_{2})$, then the third intersection point of $AB$ with $\mu({\mathbb{C}}/L)$ is $\mu(-z_{1}-z_{2})$, and then the point $A+B$ corresponds to $\mu(z_{1}+z_{2})$, so the two group laws are matched under $\mu$.12.212.212.2We are ignoring various issues of multiple roots, or one of $A$ or $B$ or $A+B$ being equal to $(0:1:0)$. These can be handled by continuity, or directly.
@@
Uniqueness is clear because the ratio of two functions with the same multiplicities of zeros and poles is holomorphic on ${\mathbb{C}}/L$ and thus constant. Existence is harder to prove – we will do it with the help of the Jacobi theta function which is a holomorphic function on $\mathbb{C}$$\times\mathcal{H}$, where $\mathcal{H}=\{\tau,{\rm~Im}\tau>0\}$ is the upper half plane.
@@
We lift $z_{k}$ to $\mathbb{C}$}. We would like to be able to do it so that $\sum_{k}m_{k}z_{k}=0$, but a priori we can only get $\sum_{k}m_{k}z_{k}\in L$, if none of the $m_{k}$ is equal to $\pm 1$. However, we can also split some $m_{k}$ into $m_{k}-1$ and $1$ and do a separate lift for $1$ to get the sum to be exactly zero. Now the elliptic function $f$ from Proposition 13.3 defines a meromorphic function on ${\mathbb{C}}/L$ with prescribed zeros and poles. (Note that we did not assume in Proposition 13.3 that $z_{k}$ are pairwise distinct modulo $L$.) ∎
@@
The torsion is well-understood, but the rank is harder. For example, it is unknown whether the rank $n$ is bounded from above. The current world record of a curve whose rank is at least $29$ is due to Elkies and Klagsbrun, achieved in 2024. The previous world record of $n\geq 28$ was achieved by Elkies in 2006, see [13].
@@
${\mathrm{i}}\sum_{n\in{\mathbb{Z}}}(-1)^{n}{\mathrm{e}}^{\pi{\mathrm{i}}((n+\frac{1}{2})^{2}\tau+(2n+1)z)}$
@@
First, we note that in the case of $X={\mathbb{C}}{\mathbb{P}}^{1}$ the degree of $K_{X_{{\mathbb{C}}{\mathbb{P}}^{1}}}$ is (2)(-2). Indeed, if we look at the differential form $dz$ which has no zeros/poles in $\mathbb{C}$}, it has a second order pole at infinity, because $dz=-\left(\frac{1}{z}\right)^{-2}\,d\left(\frac{1}{z}\right)$ near $\infty$.
@@
For $g(X)=1$, the automorphism group of $X={\mathbb{C}}/L$ is a semidirect product of ${\mathbb{C}}/L$ with a finite group, which could be ${\mathbb{Z}}/2{\mathbb{Z}}$, ${\mathbb{Z}}/4{\mathbb{Z}}$ or ${\mathbb{Z}}/6{\mathbb{Z}}$. The normal subgroup ${\mathbb{C}}/L$ is obtained by parallel transports $\varphi_{w}:z\mapsto z+w$, and the finite group comes from scalar multiplications that preserve the lattice $L$. It is usually just $\{\pm 1\}$, but it is $\{1,{\mathrm{i}},-1,{\mathrm{i}}\}$ if $L={\mathbb{Z}}+{\mathbb{Z}}{\mathrm{i}}$ and it is the group of sixth roots of $1$ for the hexagonal lattice $L={\mathbb{Z}}+{\mathbb{Z}}(\frac{1}{2}+\frac{\sqrt{3}}{2}{\mathrm{i}})$. We do not prove this, but it is not too hard to derive from what we know about elliptic curves.
@@
$2+\frac{1}{2}+\frac{2}{3}+\frac{6}{7},$
@@
$\frac{1}{2\pi{\mathrm{i}}}\oint_{|z|=\alpha}\frac{f^{\prime}(z)}{f(z)-\varepsilon}\,dz.$
It is continuous in $\varepsilon$ for $\varepsilon$ close to 0 and is equal to $\frac{1}{2\pi{\mathrm{i}}}\oint_{|z|=\alpha}\frac{f^{\prime}(z)}{f(z)}\,dz.$
@@
Exercise 3. Prove the last claim of the section. Hint: First argue that for $k\geq 5$ the right hand side is at least $\frac{1}{2}$. Then for $k=4$ the right hand is positive, so at least one of $n_{i}$ is $3$ or more, which gives at least $2+\frac{1}{2}+\frac{1}{2}+\frac{1}{2}+\frac{2}{3}=\frac{1}{6}$. If $k\leq 2$, the right hand side is negative, so we must have $k=3$, etc.
@@
$S=\{(x_{1},\ldots,x_{n})\times(y_{1}:\ldots:y_{n}),{\rm~such~that~rk}\thinspace\left(\begin{array}[]{ccc}x_{1}&\cdots&x_{n}\\ y_{1}&\cdots&y_{n}\end{array}\right)=1\}.$
@@
    $X=X_{0}\dashleftarrow\thinspace\dashrightarrow X_{1}\dashleftarrow\thinspace\dashrightarrow\cdots\dashleftarrow\thinspace\dashrightarrow X_{n-1}\dashleftarrow\thinspace\dashrightarrow X_{n}=Y$
@@
Now that we have established that $({\mathbb{C}}[x_{1},\ldots,x_{n}]/I)^{G}$ is finitely generated, let us prove the bijection between its maximal ideals and the set of orbits of $G$-action on the set of maximal ideals of ${\mathbb{C}}[x_{1},\ldots,x_{n}]/I$.16.116.116.1Recall that points in $X$ correspond to maximal ideals in $R={\mathbb{C}}[x_{1},\ldots,x_{n}]/I$ by Hilbert’s Nullestellensatz. In one direction, to any maximal ideal $m\subset R$ we associate the ideal $m^{G}=m\cap R^{G}\subset R^{G}$. The ideal $m^{G}$ is maximal because for any $x\in R^{G}\setminus m^{G}$ there exists $y\in R$ such that $xy=1\,{\rm mod}\,m$. Then the element
@@
on it. This gives a group homomorphism ${\rm SU(2)}\to{\rm SO}(3,{\mathbb{R}})$, whose kernel is $\{\pm 1\}$. Finite subgroups of ${\rm SO}(3,{\mathbb{R}})$ are classically known. They include two infinite series (rotation symmetries of a regular pyramid and a regular bipyramid) as well as rotational symmetries of the regular tetrahedron, cube/octahedron, and icosahedron/dodecahedron. This allows one to classify all finite subgroups $G\subset{\rm SL}(2,{\mathbb{C}})$.
@@
$\bullet$$\{uv=w^{2}\}\subset{\mathbb{C}}^{3}$$\tilde{S}\subset\tilde{\mathbb{C}}^{3}$
Let us see this with all the gory details, in the coordinate charts. Recall that the blowup $\tilde{\mathbb{C}}^{3}$ is a submanifold of ${\mathbb{C}}^{3}\times{\mathbb{C}}{\mathbb{P}}^{2}$, given by $\{(u,v,w),(u_{1}:v_{1}:w_{1})\}$ such that there exists $\lambda$ with $(u,v,w)=\lambda(u_{1},v_{1},w_{1})$. As a result, $\tilde{\mathbb{C}}^{3}$ is covered by three open charts, each one isomorphic to ${\mathbb{C}}^{3}$.
@@
We thus see that the proper preimage $\tilde{S}$ of $S$ is smooth. It is also instructive to look at the fibers of $\tilde{S}\to S$. Outside of $(0,0,0)$, the fibers are of course points. The preimage of $(0,0,0)$ is a subvariety of the exceptional divisor, given by $(u_{1}:v_{1}:w_{1})\in{\mathbb{C}}{\mathbb{P}}^{2}$ with $u_{1}v_{1}-w_{1}^{2}=0$, so it is a ${\mathbb{C}}{\mathbb{P}}^{1}$. For example, in the coordinate chart $w_{1}\neq 0$, we have $\frac{u}{w}=\frac{u_{1}}{w_{1}},~\frac{v}{w}=\frac{v_{1}}{w_{1}}$ so $0=\left(\frac{u}{w}\right)\left(\frac{v}{w}\right)-1$ matches $u_{1}v_{1}-w_{1}^{2}=0$ in the appropriate chart on ${\mathbb{C}}{\mathbb{P}}^{2}$.
@@
We repeat the process. Each time, the power $n$ decreases by $2$ and two new ${\mathbb{C}}{\mathbb{P}}^{1}$-s spring out. A paranoid reader can verify that under the blowup the previous ${\mathbb{C}}{\mathbb{P}}^{1}$-s separate and two new lines intersect them transversely. The last step is either unnecessary (if $n$ is odd, and we reduce to $uv=w$) or we only have one ${\mathbb{C}}{\mathbb{P}}^{1}$ (if $n$ is even and we reduce to $A_{1}$-singularity). In both cases we end up with a resolution of singularities $\tilde{S}\to S$ such that the preimage of the singular point is a chain of (n1)(n-1) transversal ${\mathbb{C}}{\mathbb{P}}^{1}$-s.
@@
with n1n-1 vertices, which is the $A_{n-1}$ Dynkin diagram.17.317.317.3There is, of course, more to this notation. Appropriately understood intersection pairing matrix of the components is up to the sign the Cartan matrix of $A_{n-1}$. We will not try to define it now, but will comment on it in Section 19.
@@
The symmetry $(x,y)\mapsto(-y,x)$ now becomes the involution $(u,v,w)\mapsto(v,u,-w)$. We can take the invariants first and then the quotient. The action of this involution on ${\mathbb{C}}^{3}$ has positive eigenspace $u+v$ and the $A_{1}$-type action on uvu-v and $w$, so the ring of invariants is generated by
@@
$2x=0,~z+(n-1)z^{n+1}y^{n-2}=0,~y+(n+1)z^{n}y^{n-1}=0,~x^{2}+yz+z^{n+1}y^{n-1}=0.$
We can multiply the second equation by $(n+1)y$ and the third one by (n1)z(n-1)z, subtract them and get $yz=0$. Since $n>2$, we use the second and third equations to deduce that both $y$ and $z$ are 0. So the only singular point in this chart is $(0,0,0)$, which corresponds to $(0:1:0)$ on the exceptional ${\mathbb{C}}{\mathbb{P}}^{2}$ of $\tilde{\mathbb{C}}^{3}\to{\mathbb{C}}^{3}$ (we are in the chart $v_{1}\neq 0$). For $n>2$, we see that the singularity at this point is of type $A_{1}$, because we can ignore higher degree terms. The intersection with the exceptional locus $v=0$ is given by $\frac{u}{v}=0$, i.e. we have the line (0::)(0:*:*) on the exceptional ${\mathbb{C}}{\mathbb{P}}^{2}$ of $\tilde{\mathbb{C}}^{3}\to{\mathbb{C}}^{3}$.
@@
$\begin{array}[]{c}{\mathbb{C}}^{2}\hskip 50.0ptS_{1}\hskip 50.0ptS_{3}\enskip\\ \searrow\qquad\swarrow\qquad\searrow\qquad\swarrow\\ {\mathbb{C}}^{2}/G_{1}\hskip 40.0ptS_{2}\quad\\ \searrow\hskip 30.0pt\swarrow\\ {\mathbb{C}}^{2}/G\end{array}$
@@
and the action of $G/G_{1}\cong{\mathbb{Z}}/2{\mathbb{Z}}$ which switches $u$ and $v$. The resolution $S_{1}$ has a chain of (2n1)(2n-1) Riemann spheres
@@
which intersect transversely, and the involution of $G/G_{1}$ switches the opposite spheres. Because we have an odd number of spheres, this involution will preserve the middle ${\mathbb{C}}{\mathbb{P}}^{1}$. It acts by switching two intersection points with two adjacent ${\mathbb{C}}{\mathbb{P}}^{1}$-s. We can call these points 0 and $\infty$ and observe that every such involution can be written in the form $z\mapsto\frac{1}{z}$, see Exercise 1. Then we have two fixed points $z=\pm 1$. Since these are isolated points, the eigenvalues of the action on the tangent space must be (1,1)(-1,-1), so the quotient will have an $A_{1}$ singularity there. Therefore, the image of the chain of (2n1)(2n-1) lines on $S_{2}$ will be $n$ lines, with the first line having two additional $A_{1}$ singularities on it. Blowing up these $A_{1}$ singularities, we get $S_{3}$ with the $(n+2)$ lines in $D_{n+2}$ configuration.
We can similarly get the $E_{6}$ singularity as the quotient of the $D_{4}$ singularity by a ${\mathbb{Z}}/3{\mathbb{Z}}$ action. Then if you resolve the $D_{4}$ singularity to get a surface with a $D_{4}$ configurations of projective lines, the group can be seen rotating the three “outside” lines while fixing the middle one. By Exercise 2, any order three automorphism of ${\mathbb{C}}{\mathbb{P}}^{1}$ can be written in the form $z\mapsto{\mathrm{e}}^{2\pi{\mathrm{i}}/3}z$, so it will have two fixed points 0 and $\infty$. It is then possible to compute that the quotient will be the $A_{2}$ singularity, so we get $S_{1}\to S$ with $E_{6}$ singularity replaced by two lines with two $A_{2}$ singular points on one of them.
@@
Exercise 1. Prove that every involution of ${\mathbb{C}}{\mathbb{P}}^{1}$ that sends 0 to $\infty$ can be written in the form $z\mapsto\frac{c}{z}$ for some nonzero constant $c$. Then argue that by scaling $z$ we can rewrite it as $z\mapsto\frac{1}{z}$.
@@
If Y\,Y and $Z$ intersect transversely at $k$ points, then $Y\cdot Z=k$ (in particular, $Y\cdot Z=0$ if $Y\cap Z=\emptyset$). More generally, if $Y\cap Z$ is a finite set, then $Y\cdot Z$ is a sum of local contributions, each of which is positive. However, the definition makes sense even if $Y\cap Z$ is positive-dimensional, or even $Y\subseteq Z$! In these cases, we just know that $Y\cdot Z$ is integer but can make no claims about the sign of it. Philosophically, we can deform $Y$ as a real cycle to some $Y^{\prime}$ with the same homology classes $Y^{\prime}]=[Y]$ so that $Y^{\prime}$ intersects $Z$ transversely, and we add up $(\pm 1)$ at each intersection point, depending on the orientations.
@@
The key observation is that $L\sim L^{\prime}+E$, in the sense that $L]=[L^{\prime}]+[E]$ in $H_{2}(X,{\mathbb{Z}})$. Indeed, more generally there is a map of Weil divisors to $H_{2}(X,{\mathbb{Z}})$ which passes through the class group. Another way of saying the same thing is that a divisor of a rational function is always homologically trivial. And there is a rational function (the ratio of the equations of the equations of the two lines) which has simple zeros at $E$ and $L^{\prime}$ and poles at LL.19.319.319.3In this particular case can think of the family of smooth preimages of lines in ${\mathbb{C}}{\mathbb{P}}^{2}$ degenerating into a singular curve $E\cup L^{\prime}$. We also have
@@
$C_{1}^{\prime}]+[E]+[C_{2}^{\prime}]-[D]$
where $D$ is the preimage of the conic at infinity $\{uv-w^{2}=t=0\}$. Therefore, we have $E]=[D]-[C_{1}^{\prime}]-[C_{2}^{\prime}]$ in $H_{2}(X,{\mathbb{Z}})$ and
@@
Exercise 3. Let $X$ be a blowup of ${\mathbb{C}}{\mathbb{P}}^{2}$ in $5$ points $p_{1},\ldots,p_{5}$ in general position. Prove that the proper preimage of any line through two of these points has self-intersection (1)(-1). Prove that the proper preimage of the unique conic that passes through all five points also has self-intersection (1)(-1).
@@
One can show that every Weil divisor on ${\mathbb{C}}^{2}$ can be given as a divisor of a rational function. Therefore, the class group of $H_{n}$ is generated by $S_{0}]$ and $F_{0}]$. Let us find their intersection numbers.
@@
The rational function $\frac{x_{0}}{x_{1}}$ on ${\mathbb{C}}{\mathbb{P}}^{1}$ gives a rational function on $H_{n}$ with divisor $F_{0}]-[F_{\infty}]$. We thus get
@@
$S_{0}]-[S_{-n}]+n[F_{0}].$ (20.2)
This would imply that $S_{0}]=[S_{-n}]-n[F_{0}]$ in ${\rm Cl}(H_{n})$, so
@@
has divisor $S_{0}]-[S_{-n}]$ in the first chart and $S_{0}]-[S_{-n}]+n[F_{0}]$ in the second. ∎